# On the Geometric Coherence of Global Aggregation in Federated GNN

### v1

@@ 2.4  Mismatch Induced by Linear Aggregation
However, the update $\Delta\theta_{k}$ was optimized with respect to $\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}$, not $\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}}$. Unless the Jacobians $\{\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}\}$ share aligned dominant eigenspaces or consistent sensitivity profiles, Equation (16) does not correspond to a coherent operator update for any $\mathcal{G}$.Geometrically, FedAvg performs linear combination of vectors $\{\Delta\theta_{k}\}$ that originate from incompatible operator tangent spaces, implicitly projecting them onto a shared parameter space without preserving their induced transformation geometry.
2.5  Demonstration of Destructive Interference
Consider two clients $i$ and $j$ whose dominant operator perturbations satisfy
@@ 2.5  Demonstration of Destructive Interference
$\langle\mathcal{J}_{\theta}^{\mathcal{G}_{i}}\Delta\theta_{i},\,\mathcal{J}_{\theta}^{\mathcal{G}_{j}}\Delta\theta_{j}\rangle<0.$ (17)
In this case, aggregation induces partial cancellation in the operator space, even if $\|\Delta\theta_{i}\|$ and $\|\Delta\theta_{j}\|$ are individually large. Such cancellation reduces the effective rank or sensitivity of the global operator, leading to contraction of message-passing dynamics.This phenomenon is not captured by parameter norms or loss values, as it arises from misalignment in the induced operator geometry.In the Next section, we discuss Given heterogeneous graph distributions $\{\mathcal{D}_{k}\}$ and locally optimized updates $\{\Delta\theta_{k}\}$, standard federated aggregation constructs a global parameter update that does not, in general, correspond to a geometrically coherent perturbation of the induced message-passing operator.
2.6  Geometric Regulation of Federated GNN Aggregation
Revisiting the Source of Mismatch:
From Section III, local updates $\Delta\theta_{k}$ induce graph-dependent operator perturbations
$\Delta\mathcal{T}_{k}\;=\;\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\Delta\theta_{k},$ (18)
which generally lie in incompatible tangent spaces due to heterogeneity in $\mathcal{G}_{k}$.
We emphasize that the Jacobian-based operator perturbation view in Eqs. (14)–(18) is intended as a first-order conceptual lens for understanding why parameter-space averaging may distort message-passing semantics under heterogeneous client graphs. In practice, the server does not have access to the true client-specific Jacobians $J_{\theta}^{G_{k}}$, nor does it explicitly construct operators in closed form. Rather, the key insight is that local updates are graph-conditioned: the same parameter displacement $\Delta\theta_{k}$ can induce qualitatively different propagation behavior depending on the underlying adjacency structure, feature statistics, and spectral regime of each client. Thus, geometric incompatibility should be interpreted as a mismatch in the directional and sensitivity characteristics of client updates, rather than as an exact computation in an abstract operator manifold. This motivates the use of lightweight server-side proxy measures (such as layerwise directional coherence) to regulate aggregation and preserve stable global message-passing dynamics.Standard aggregation performs linear combination in parameter space, implicitly assuming
$\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\approx\mathcal{J}_{\theta}^{\mathcal{G}_{k^{\prime}}},$ (19)
an assumption that does not hold in federated GNNs.The core objective is therefore not to align parameters directly, but to ensure that the induced operator perturbations remain geometrically compatible under aggregation.
Operator-Space Reformulation of Aggregation:
Let $\mathcal{O}$ denote the space of admissible message-passing operator perturbations. Each client update induces an element $\Delta\mathcal{T}_{k}\in\mathcal{O}$. The server observes $\{\Delta\theta_{k}\}$ but does not have access to $\mathcal{G}_{k}$. However, the effect of each update can be characterized through its action on the current global operator.
We therefore define a server-side mapping
$\Psi:\Delta\theta_{k}\mapsto\widehat{\Delta\mathcal{T}}_{k},$ (20)
where $\widehat{\Delta\mathcal{T}}_{k}$ is a proxy representation of the induced operator perturbation, constructed without access to client data.
Geometric Admissibility Constraints:
We characterize the geometric coherence of operator perturbations through three structural properties: Directional Consistency:Let $\langle\cdot,\cdot\rangle_{\mathcal{O}}$ denote an inner product in operator space. An update $\widehat{\Delta\mathcal{T}}_{k}$ is directionally consistent with the current global operator if
$\langle\widehat{\Delta\mathcal{T}}_{k},\Delta\mathcal{T}^{\text{ref}}\rangle_{\mathcal{O}}\geq 0,$ (21)
where $\Delta\mathcal{T}^{\text{ref}}$ denotes a reference global direction (e.g., a running estimate of admissible operator change). This constraint prevents semantic inversion of message passing under aggregation. Subspace Compatibility:Let $\mathcal{S}\subset\mathcal{O}$ denote a low-dimensional subspace spanned by the dominant propagation modes of the global operator. Each update is projected as
$\Pi_{\mathcal{S}}(\widehat{\Delta\mathcal{T}}_{k}),$ (22)
and aggregation is restricted to this subspace. This preserves diversity of admissible propagation modes while preventing collapse into degenerate directions. Sensitivity Stability:Let $\|\cdot\|_{\mathcal{O}}$ denote a norm capturing sensitivity of the operator to graph perturbations. We impose boundedness:
@@ 2.6  Geometric Regulation of Federated GNN Aggregation
$\|\widehat{\Delta\mathcal{T}}_{k}\|_{\mathcal{O}}\leq\epsilon,$ (23)
preventing excessive contraction or amplification of message passing.
Geometrically Regulated Aggregation
Under these constraints, server-side aggregation is reformulated as
$\Delta\theta^{\text{global}}=\sum_{k=1}^{K}w_{k}\,\mathcal{R}(\Delta\theta_{k}),$ (24)
where $\mathcal{R}(\cdot)$ denotes a geometric regulation operator that enforces directional consistency, subspace compatibility, and sensitivity stability through the proxy operator representation $\widehat{\Delta\mathcal{T}}_{k}$. Importantly, $\mathcal{R}$ operates solely on transmitted updates and does not require access to client graphs or data.
Preservation of Operator Coherence
Let the regulated operator perturbation be
$\Delta\mathcal{T}^{\text{reg}}=\sum_{k=1}^{K}w_{k}\,\widehat{\Delta\mathcal{T}}_{k}^{\text{reg}}.$ (25)
By construction,
$\Delta\mathcal{T}^{\text{reg}}\in\mathcal{S},\quad\langle\Delta\mathcal{T}^{\text{reg}},\Delta\mathcal{T}^{\text{ref}}\rangle_{\mathcal{O}}\geq 0,\quad\|\Delta\mathcal{T}^{\text{reg}}\|_{\mathcal{O}}\leq\epsilon.$ (26)
Thus, the updated global model
$\theta^{(t+1)}=\theta^{(t)}+\Delta\theta^{\text{global}}$ (27)
induces a message-passing operator whose geometry remains coherent across rounds. Hence,we draw inferences on the effects imposed on Iterative Federated Dynamics.The regulated global model preserves operator geometry.The broadcast model serves as a geometrically aligned initialization for subsequent local optimization.Therefore,Local updates act as refinements within an admissible operator neighborhood,rather than corrective responses to geometric distortion.This breaks the destructive feedback loop identified in previous comprehension and stabilizes federated GNN training under heterogeneous graph distributions.The proposed regulation neither modifies local training nor does it alter the task objective.It does not assume homogeneity across clients as well.Instead, it corrects a structural mismatch by enforcing geometric coherence at the only location where heterogeneous updates interact—the global server.
    ethod
Further, we describe the concrete instantiation of the proposed Global Geometric Reference Structure (GGRS) and its integration into the federated aggregation pipeline. We emphasize that GGRS is a server-side mechanism that operates exclusively on client model updates and does not require access to client data, graph topology, or labels. In FL, the global server is the only entity that simultaneously observes updates from multiple clients. Local clients optimize their models independently and have no visibility into the distribution or interaction of updates produced by other clients. Consequently, any geometric incompatibility arising from the interaction of heterogeneous updates can only be detected and corrected at the server.Furthermore, the global model produced by aggregation serves both as the predictor used for inference and as the initialization for subsequent local optimization rounds. As shown in Section IV, geometric distortion introduced at this stage propagates forward into prediction behavior and backward into local training dynamics. Therefore, preserving geometric coherence at the global server is both necessary and sufficient to stabilize the federated GNN system. GGRS operates by constructing and maintaining a global geometric reference in the space of induced message-passing operator perturbations. At each communication round, this reference serves as an anchor against which incoming client updates are evaluated and regulated.As shown in 3,the construction consists of three components:A proxy representation of operator perturbations induced by client updates;A dynamically maintained reference geometry capturing admissible global propagation behavior and a regulation operator that aligns client updates with the reference prior to aggregation.In cross-domain federation, GGRS regulates and aggregates only the shared encoder parameters, while task-specific classifier heads remain local.
Proxy Representation of Operator Perturbations: Let $\theta^{(t)}$ denote the global model at round $t$, and let $\Delta\theta_{k}^{(t)}$ be the update received from client $k$. As the server does not have access to client graphs, it cannot compute the true operator perturbation $\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\Delta\theta_{k}^{(t)}$. Instead, GGRS constructs a proxy operator representation based on the effect of the update on the current global model.Specifically, the server evaluates the directional effect of $\Delta\theta_{k}^{(t)}$ on the global model parameters and maps it into a low-dimensional geometric summary:
@@
$\mathbf{z}_{k}^{(t)}=\Psi\left(\Delta\theta_{k}^{(t)},\theta^{(t)}\right),$ (28)
where $\Psi(\cdot)$ is a deterministic mapping that captures dominant update directions and relative scaling across layers. In practice, $\Psi$ may be instantiated using layer-wise normalized update projections or low-rank approximations of parameter perturbations. The resulting vector $\mathbf{z}_{k}^{(t)}$ serves as a surrogate for the induced operator perturbation and enables geometric comparison across clients.
Global Geometric Reference Construction: GGRS maintains a reference vector $\mathbf{r}^{(t)}$ that summarizes the admissible direction of global message-passing evolution. This reference is updated incrementally across communication rounds as:
$\mathbf{r}^{(t)}=\alpha\mathbf{r}^{(t-1)}+(1-\alpha)\sum_{k=1}^{K}w_{k}\mathbf{z}_{k}^{(t)},$ (29)
where $\alpha\in[0,1)$ controls the temporal smoothing of the reference and $w_{k}$ are aggregation weights. The reference $\mathbf{r}^{(t)}$ does not represent an optimal update direction for any individual client. Rather, it encodes a geometrically admissible envelope within which client updates can be combined without inducing operator degeneration.
Geometric Regulation of Client Updates: Given the proxy representation $\mathbf{z}_{k}^{(t)}$ and the reference $\mathbf{r}^{(t)}$, GGRS regulates each client update via three operations: 1) Directional Alignment:Updates whose proxy representations exhibit negative alignment with the reference are reflected or downweighted:
$\mathbf{z}_{k}^{\prime}=\begin{cases}\mathbf{z}_{k}^{(t)},&\langle\mathbf{z}_{k}^{(t)},\mathbf{r}^{(t)}\rangle\geq 0,\\ \beta\mathbf{z}_{k}^{(t)},&\text{otherwise},\end{cases}$ (30)
with $\beta\in[0,1)$.Subspace Projection: To prevent collapse into degenerate propagation modes, regulated updates are projected onto a low-dimensional subspace $\mathcal{S}^{(t)}$ spanned by the dominant directions of recent proxy updates:
$\mathbf{z}_{k}^{\prime\prime}=\Pi_{\mathcal{S}^{(t)}}(\mathbf{z}_{k}^{\prime}).$ (31)
Sensitivity Normalization: Finally, updates are rescaled to enforce bounded sensitivity:
$\widehat{\mathbf{z}}_{k}^{(t)}=\frac{\mathbf{z}_{k}^{\prime\prime}}{\max(1,\|\mathbf{z}_{k}^{\prime\prime}\|/\epsilon)}.$ (32)
The regulated proxy $\widehat{\mathbf{z}}_{k}^{(t)}$ is then mapped back to parameter space by applying the same regulation coefficients to $\Delta\theta_{k}^{(t)}$.
GGRS-Enabled Aggregation: The regulated global update is computed as:
$\Delta\theta^{(t)}=\sum_{k=1}^{K}w_{k}\mathcal{R}\left(\Delta\theta_{k}^{(t)}\right),$ (33)
where $\mathcal{R}(\cdot)$ denotes the regulation induced by the proxy operations described above. The global model is updated as:
$\theta^{(t+1)}=\theta^{(t)}+\Delta\theta^{(t)}.$ (34)
We propose a GGRS Algorithm as stated below, and the 3 visualizes the overall GGRS Construction.
Algorithm 1 Federated GNN Training with Global Geometric Regulation (GGRS)
0: Initial global shared encoder $\theta^{(0)}$, number of rounds $T$, clients $\{1,\dots,K\}$, proxy mapping $\Psi(\cdot)$, regulation operator $\mathcal{R}(\cdot)$
1: Initialize geometric reference $\mathbf{r}^{(0)}\leftarrow\mathbf{0}$
2:for $t=0,1,\dots,T-1$ do
3:  Server broadcasts shared encoder $\theta^{(t)}$ to all clients
4:  for all clients $k$ in parallel do
5:   Initialize local encoder with $\theta^{(t)}$ (heads remain local)
6:   Train local GNN on private graph $\mathcal{G}_{k}$
7:   Compute encoder update $\Delta\theta_{k}^{(t)}=\theta_{k}^{(t)}-\theta^{(t)}$
8:   Send $\Delta\theta_{k}^{(t)}$ to server
9:  end for
10:  for all received updates $\Delta\theta_{k}^{(t)}$ do

### v2

@@ 2.3  Operator Perturbations and Geometric Incompatibility
Because each $\Delta\theta_{k}$ was optimized under $\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}$ rather than $\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}}$, the sum in (7) conflates vectors from geometrically incompatible tangent spaces. Unless $\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}\approx\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k^{\prime}}}$ for all $k,k^{\prime}$—an assumption that fails whenever clients differ substantially in graph structure [7, 5]—the resulting global operator lacks coherent directional structure.
Destructive interference.  Consider two clients $i,j$ satisfying
$\bigl\langle\mathcal{J}_{\theta}^{\mathcal{G}_{i}}\Delta\theta_{i},\;\mathcal{J}_{\theta}^{\mathcal{G}_{j}}\Delta\theta_{j}\bigr\rangle<0.$ (8)
Their combined contribution to $\Delta\mathcal{T}^{\mathcal{G}}$ partially cancels in operator space, even when $\|\Delta\theta_{i}\|$ and $\|\Delta\theta_{j}\|$ are each large. This cancellation reduces the effective rank and sensitivity of $\Delta\mathcal{T}^{\mathcal{G}}$, contracting message-passing dynamics in a manner invisible to parameter norms or training loss.
Training feedback amplification.  The distorted $\theta^{(t+1)}$ is broadcast to all clients as the initialization for round t+1t{+}1. Each client must now apply a corrective update to recover operator geometry suited to $\mathcal{G}_{k}$, but these corrective updates are themselves mutually misaligned, amplifying interference at the subsequent aggregation step. This feedback loop constitutes the central failure mode in heterogeneous federated GNN training.
This leads us to our problem statement: given a federated system with heterogeneous graph distributions $\{\mathcal{D}_{k}\}_{k=1}^{K}$, we are to design a server-side aggregation mechanism that preserves the geometric coherence of the induced message-passing operator—specifically, directional alignment, low-dimensional subspace structure, and sensitivity stability—across communication rounds, without access to client graphs or data.
    ormal Derivation of Geometric Regulation in Federated GNN Aggregation
3.1  Federated Aggregation as Tangent-Space Averaging
Recall from (6) that local update $\Delta\theta_{k}$ induces an operator perturbation $\Delta\mathcal{T}_{k}=\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\Delta\theta_{k}$. The FedAvg global perturbation is
$\Delta\mathcal{T}^{\text{FedAvg}}=\sum_{k=1}^{K}w_{k}\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\Delta\theta_{k}.$ (9)
Equation (9) is a weighted sum of vectors from distinct tangent spaces $\{\mathcal{T}_{\theta}^{\mathcal{G}_{k}}\}$. Its deficiency can be quantified via the cosine misalignment:
@@ 3.1  Federated Aggregation as Tangent-Space Averaging
$\cos\!\angle\bigl(\Delta\mathcal{T}_{i},\,\Delta\mathcal{T}_{j}\bigr)=\frac{\langle\mathcal{J}_{\theta}^{\mathcal{G}_{i}}\Delta\theta_{i},\;\mathcal{J}_{\theta}^{\mathcal{G}_{j}}\Delta\theta_{j}\rangle}{\|\mathcal{J}_{\theta}^{\mathcal{G}_{i}}\Delta\theta_{i}\|\;\|\mathcal{J}_{\theta}^{\mathcal{G}_{j}}\Delta\theta_{j}\|}.$ (10)
When (10) is negative for a significant fraction of client pairs, the aggregated perturbation suffers destructive interference as formalized in (8).
3.2  The Data-Free Proxy Map $\Psi$
The fundamental obstacle is that the server has no access to $\{\mathcal{G}_{k}\}$ and therefore cannot compute $\Delta\mathcal{T}_{k}$ exactly. We derive a data-free proxy $\widehat{\Delta\mathcal{T}}_{k}$ that captures the directional signature of $\Delta\mathcal{T}_{k}$ from $\Delta\theta_{k}$ alone.
Step 1: Factoring the Jacobian.
Write the composite Jacobian as
$\mathcal{J}_{\theta}^{\mathcal{G}_{k}}=\mathbf{P}_{k}\,\mathbf{Q},$ (11)
where $\mathbf{Q}\in\mathbb{R}^{p\times p}$ is a graph-independent component capturing global parameter sensitivity (e.g., $\mathbf{Q}=\mathbf{I}_{p}$ as the first-order approximation, or the empirical Fisher information matrix averaged across clients), and $\mathbf{P}_{k}$ encodes the graph-specific modulation of propagation. Under moderate heterogeneity the dominant left singular subspace of $\mathbf{P}_{k}$ is approximately shared across clients [7, 15], motivating the approximation
$\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\;\approx\;\mathbf{Q},\quad\forall k\in[K].$ (12)
Step 2: Deriving the proxy direction.
Under approximation (12), the induced operator perturbation reduces to
$\Delta\mathcal{T}_{k}\;\approx\;\mathbf{Q}\,\Delta\theta_{k}.$ (13)
Taking $\mathbf{Q}=\mathbf{I}_{p}$ as the canonical reference metric (corresponding to the limit of uniform parameter sensitivity), the directional component of the operator perturbation is characterized entirely by the unit vector
$\mathbf{z}_{k}\;=\;\frac{\Delta\theta_{k}}{\|\Delta\theta_{k}\|_{2}}\;\in\;\mathbb{S}^{p-1},$ (14)
which belongs to the unit hypersphere in parameter space.
Step 3: Formal definition of $\Psi$.
We define the data-free proxy map
$\Psi:\;\Delta\theta_{k}\;\longmapsto\;\widehat{\Delta\mathcal{T}}_{k}\;\triangleq\;\frac{\Delta\theta_{k}}{\|\Delta\theta_{k}\|_{2}},$ (15)
mapping each transmitted update to its unit-normalized direction in parameter space. $\Psi$ is well-defined for $\Delta\theta_{k}\neq\mathbf{0}$ and requires only the transmitted update vector; it does not access $\mathcal{G}_{k}$.
Remark 1.
The identity approximation $\mathbf{Q}=\mathbf{I}_{p}$ is used throughout for tractability. When a global estimate of the Fisher information matrix $\hat{\mathbf{F}}$ is available (e.g., from a proxy dataset), one may use $\Psi(\Delta\theta_{k})=\hat{\mathbf{F}}^{1/2}\Delta\theta_{k}/\|\hat{\mathbf{F}}^{1/2}\Delta\theta_{k}\|_{2}$, which weights parameter directions by their sensitivity and provides a tighter proxy. The identity case is equivalent to treating all parameter dimensions as equally informative, which is a standard assumption in parameter-space federated averaging.
3.3  Reference Direction and Geometric Admissibility
Constructing the reference operator direction.
Given the proxy vectors $\{\mathbf{z}_{k}^{(t)}\}=\{\Psi(\Delta\theta_{k}^{(t)})\}$, we maintain an exponential moving average (EMA) reference direction tracking the dominant aggregated perturbation over time:
$\mathbf{r}^{(t)}=\alpha\,\mathbf{r}^{(t-1)}+(1-\alpha)\sum_{k=1}^{K}w_{k}\,\mathbf{z}_{k}^{(t)},\quad\mathbf{r}^{(0)}=\mathbf{0},$ (16)
with decay parameter $\alpha\in(0,1)$. After normalization, the unit reference vector is
$\hat{\mathbf{r}}^{(t)}=\frac{\mathbf{r}^{(t)}}{\|\mathbf{r}^{(t)}\|_{2}+\epsilon_{0}},\quad\epsilon_{0}>0\text{ (numerical stabilizer)}.$ (17)
The EMA in (16) serves as a server-side estimate of the admissible operator perturbation direction: it accumulates the weighted consensus of past proxy directions, giving more weight to recent rounds via $\alpha$.
Three geometric admissibility constraints.
We characterize coherence of an operator perturbation $\widehat{\Delta\mathcal{T}}_{k}$ relative to $\hat{\mathbf{r}}^{(t)}$ through three structural conditions.
C1 (Directional Consistency).  The cosine alignment of client $k$’s proxy with the reference direction is
$\gamma_{k}^{(t)}\;=\;\bigl\langle\mathbf{z}_{k}^{(t)},\;\hat{\mathbf{r}}^{(t)}\bigr\rangle\;=\;\cos\angle\!\bigl(\Delta\theta_{k}^{(t)},\,\mathbf{r}^{(t)}\bigr).$ (18)
A client is deemed directionally consistent if $\gamma_{k}^{(t)}\geq 0$. When $\gamma_{k}^{(t)}<0$, the induced operator perturbation is antidirectional to the global consensus—its inclusion in aggregation constitutes destructive interference as in (8).
C2 (Subspace Compatibility).  Let $\mathcal{B}^{(t)}=\{\mathbf{z}_{k}^{(s)}:s\in\mathcal{W}^{(t)},k\in[K]\}$ denote the sliding buffer of proxy directions over a recent window $\mathcal{W}^{(t)}$, and form the matrix $\mathbf{Z}^{(t)}\in\mathbb{R}^{p\times|\mathcal{B}^{(t)}|}$. The admissible subspace is defined as the column space of the leading $q$ right singular vectors:
$\mathbf{Z}^{(t)}=\mathbf{U}^{(t)}\boldsymbol{\Sigma}^{(t)}{\mathbf{V}^{(t)}}^{\top},\qquad\mathcal{S}^{(t)}=\operatorname{col}\bigl(\mathbf{U}^{(t)}_{:,1:q}\bigr)\;\subset\;\mathbb{R}^{p},$ (19)
capturing the $q$ dominant modes of variation in historical proxy directions. The orthogonal projector onto $\mathcal{S}^{(t)}$ is
$\boldsymbol{\Pi}_{\mathcal{S}^{(t)}}=\mathbf{U}^{(t)}_{:,1:q}\,{\mathbf{U}^{(t)}_{:,1:q}}{}^{\top}$. (20)
A proxy direction satisfies subspace compatibility if $\boldsymbol{\Pi}_{\mathcal{S}^{(t)}}\mathbf{z}_{k}^{(t)}$ retains significant energy, i.e., components outside $\mathcal{S}^{(t)}$ are treated as noise arising from client-specific graph artifacts rather than globally informative signal.
C3 (Sensitivity Stability).  We impose an $\ell_{2}$-norm bound on the regulated proxy:
$\|\widehat{\Delta\mathcal{T}}_{k}^{\,\text{reg}}\|_{2}\;\leq\;\varepsilon,$ (21)
preventing any single client from dominating the aggregated operator perturbation through an abnormally large update magnitude. This bound also limits the sensitivity of the global operator to adversarial or outlier clients.
3.4  Derivation of the Geometric Regulation Operator $\mathcal{R}$
Given the three admissibility conditions C1–C3, we now derive $\mathcal{R}:\mathbb{R}^{p}\to\mathbb{R}^{p}$ as the composition of three sequentially applied operations.
Step 1: Directional attenuation (enforcing C1).
Define the attenuation factor
$\beta_{k}^{(t)}=\begin{cases}1,&\text{if }\gamma_{k}^{(t)}\geq 0\text{ or }t\leq w,\\[4.0pt] \beta\,\bigl(1-\gamma_{k}^{(t)}\bigr),&\text{if }\gamma_{k}^{(t)}<0\text{ and }t>w,\end{cases}$ (22)
where $\beta\in(0,1)$ is the attenuation strength hyperparameter and $w$ is a warm-up period allowing $\hat{\mathbf{r}}^{(t)}$ to stabilize. Note that $\gamma_{k}^{(t)}<0$ implies $1-\gamma_{k}^{(t)}>1$, so $\beta_{k}^{(t)}\in(0,2\beta)$; in practice we additionally clamp $\beta_{k}^{(t)}\leq\beta_{\max}<1$ to avoid full suppression. The attenuated proxy is
$\tilde{\mathbf{z}}_{k}^{(t)}=\beta_{k}^{(t)}\;\mathbf{z}_{k}^{(t)}.$ (23)
By construction, $\langle\tilde{\mathbf{z}}_{k}^{(t)},\hat{\mathbf{r}}^{(t)}\rangle=\beta_{k}^{(t)}\gamma_{k}^{(t)}$. When $\gamma_{k}^{(t)}<0$, the attenuation strictly reduces the magnitude of the antidirectional contribution without forcing it to zero, preserving a fraction of the client’s update to avoid complete exclusion.
Step 2: Subspace projection (enforcing C2).
Project the attenuated proxy onto the admissible subspace:
$\hat{\mathbf{z}}_{k}^{(t)}=\boldsymbol{\Pi}_{\mathcal{S}^{(t)}}\,\tilde{\mathbf{z}}_{k}^{(t)}=\mathbf{U}^{(t)}_{:,1:q}\,{\mathbf{U}^{(t)}_{:,1:q}}{}^{\top}\tilde{\mathbf{z}}_{k}^{(t)}.$ (24)
This discards components of $\Delta\theta_{k}$ that lie outside the dominant $q$-dimensional subspace of recent consensus directions, removing client-idiosyncratic noise orthogonal to the historical aggregation manifold.
Step 3: Sensitivity clipping (enforcing C3).
The regulated proxy is obtained by rescaling the projected direction to satisfy the sensitivity bound:
@@ 3.4  Derivation of the Geometric Regulation Operator $\mathcal{R}$
$\widehat{\Delta\mathcal{T}}_{k}^{\,\text{reg}}=\min\!\left(1,\;\frac{\varepsilon}{\|\hat{\mathbf{z}}_{k}^{(t)}\|_{2}+\epsilon_{0}}\right)\hat{\mathbf{z}}_{k}^{(t)}.$ (25)
Clipping is norm-preserving in direction and acts only when $\|\hat{\mathbf{z}}_{k}^{(t)}\|_{2}>\varepsilon$.
Composition: the regulation operator.
Combining Steps 1–3, the geometric regulation operator $\mathcal{R}$ is defined as the composition
@@
$\mathcal{R}\!\left(\Delta\theta_{k}^{(t)}\right)\;\triangleq\;\min\!\left(1,\frac{\varepsilon}{\|\hat{\mathbf{z}}_{k}^{(t)}\|_{2}+\epsilon_{0}}\right)\boldsymbol{\Pi}_{\mathcal{S}^{(t)}}\!\left[\beta_{k}^{(t)}\;\Psi\!\left(\Delta\theta_{k}^{(t)}\right)\right],$ (26)
which operates solely on the transmitted $\Delta\theta_{k}^{(t)}$ and the server-maintained quantities $\hat{\mathbf{r}}^{(t)}$ and $\boldsymbol{\Pi}_{\mathcal{S}^{(t)}}$. The regulated global update is
$\Delta\theta^{\text{global}}=\sum_{k=1}^{K}w_{k}\;\mathcal{R}\!\left(\Delta\theta_{k}^{(t)}\right).$ (27)
3.5  Theoretical Properties of $\mathcal{R}$
We now show that the global operator perturbation produced by (27) satisfies all three admissibility conditions by construction.
Proposition 1 (Geometric Admissibility of Regulated Aggregation).
Let $\Delta\theta^{\mathrm{global}}$ be given by (27) with $\mathcal{R}$ as in (26). Define the regulated global proxy $\Delta\mathcal{T}^{\mathrm{reg}}=\sum_{k=1}^{K}w_{k}\widehat{\Delta\mathcal{T}}_{k}^{\,\mathrm{reg}}$. Then:
  1. (i)
[…]
  2. (ii)
[…]
  3. (iii)
[…]
Proof.
(i) Directional consistency. Decompose the weighted sum:
$\langle\Delta\mathcal{T}^{\mathrm{reg}},\hat{\mathbf{r}}^{(t)}\rangle$ $=\sum_{k=1}^{K}w_{k}\langle\widehat{\Delta\mathcal{T}}_{k}^{\,\mathrm{reg}},\,\hat{\mathbf{r}}^{(t)}\rangle$
$=\sum_{k=1}^{K}w_{k}\,c_{k}\,\langle\boldsymbol{\Pi}_{\mathcal{S}^{(t)}}\tilde{\mathbf{z}}_{k}^{(t)},\;\hat{\mathbf{r}}^{(t)}\rangle,$ (28)
where $c_{k}=\min(1,\varepsilon/\|\hat{\mathbf{z}}_{k}^{(t)}\|_{2})\geq 0$. Note that $\hat{\mathbf{r}}^{(t)}\in\mathcal{S}^{(t)}$ by construction: since $\mathcal{S}^{(t)}$ is spanned by the leading singular vectors of $\mathbf{Z}^{(t)}$, which includes the aggregated proxy directions $\mathbf{z}_{k}^{(s)}$ for recent rounds, and $\hat{\mathbf{r}}^{(t)}$ is a convex combination of those directions, we have $\hat{\mathbf{r}}^{(t)}\in\mathcal{S}^{(t)}$. Therefore,
$\langle\boldsymbol{\Pi}_{\mathcal{S}^{(t)}}\tilde{\mathbf{z}}_{k}^{(t)},\;\hat{\mathbf{r}}^{(t)}\rangle=\langle\tilde{\mathbf{z}}_{k}^{(t)},\;\boldsymbol{\Pi}_{\mathcal{S}^{(t)}}^{\top}\hat{\mathbf{r}}^{(t)}\rangle=\langle\tilde{\mathbf{z}}_{k}^{(t)},\;\hat{\mathbf{r}}^{(t)}\rangle=\beta_{k}^{(t)}\,\gamma_{k}^{(t)},$ (29)
where we used $\boldsymbol{\Pi}_{\mathcal{S}}^{\top}=\boldsymbol{\Pi}_{\mathcal{S}}$ (orthogonal projector) and $\boldsymbol{\Pi}_{\mathcal{S}}\hat{\mathbf{r}}=\hat{\mathbf{r}}$. From the definition of $\beta_{k}^{(t)}$ in (22): if $\gamma_{k}^{(t)}\geq 0$, then $\beta_{k}^{(t)}\gamma_{k}^{(t)}\geq 0$; if $\gamma_{k}^{(t)}<0$, then $\beta_{k}^{(t)}=\beta(1-\gamma_{k}^{(t)})>0$ and $\gamma_{k}^{(t)}<0$, so $\beta_{k}^{(t)}\gamma_{k}^{(t)}<0$, but with $|\beta_{k}^{(t)}\gamma_{k}^{(t)}|<\beta|\gamma_{k}^{(t)}|$. Summing over $k$ with weights $w_{k}$, the positive contributions (from clients with $\gamma_{k}\geq 0$) dominate because the attenuation $\beta<1$ suppresses antidirectional contributions. In particular, for the dominant consensus client $k^{*}=\arg\max_{k}w_{k}\gamma_{k}$, the signed sum $\sum_{k}w_{k}\beta_{k}^{(t)}\gamma_{k}^{(t)}\geq 0$ when the weighted mean alignment of the unattenuated group exceeds $\beta\,|\sum_{k:\gamma_{k}<0}w_{k}\gamma_{k}|$, which holds by design of $\beta$.
(ii) Subspace compatibility. From (24), each $\hat{\mathbf{z}}_{k}^{(t)}=\boldsymbol{\Pi}_{\mathcal{S}^{(t)}}\tilde{\mathbf{z}}_{k}^{(t)}\in\mathcal{S}^{(t)}$. Since $\mathcal{S}^{(t)}$ is a linear subspace and $\Delta\mathcal{T}^{\mathrm{reg}}$ is a weighted sum of elements of $\mathcal{S}^{(t)}$, $\Delta\mathcal{T}^{\mathrm{reg}}\in\mathcal{S}^{(t)}$.
(iii) Sensitivity stability. From (25), $\|\widehat{\Delta\mathcal{T}}_{k}^{\,\mathrm{reg}}\|_{2}=\min\!\bigl(\|\hat{\mathbf{z}}_{k}^{(t)}\|_{2},\,\varepsilon\bigr)\leq\varepsilon$. ∎
3.6  The GGRS Algorithm
Algorithm 1 summarizes the complete server-side geometric regulation procedure. Equations (15)–(27) define the transformation applied to each received update before aggregation.
Algorithm 1 GGRS: Geometric Regulation of Federated GNN Aggregation
1:Global model $\theta^{(0)}$, weights $\{w_{k}\}$, hyperparameters $\alpha,\beta,\varepsilon,q,w$
2:Initialize $\mathbf{r}^{(0)}\leftarrow\mathbf{0}$,  $\mathcal{B}\leftarrow\emptyset$
3:for $t=1,2,\dots,T$ do
4:  Broadcast $\theta^{(t)}$ to all clients
5:  Receive $\{\Delta\theta_{k}^{(t)}\}_{k=1}^{K}$ from clients
6:  for all $k\in[K]$ do